OpenStudy (anonymous):

A computer ink cartridge has a life of X hours. The variable X is modelled by the probability density function \[f(x)=\left(\begin{matrix}kx^{-2} \\ 0\end{matrix}\right)\]x ≥ 0 otherwise (a) Find K

OpenStudy (anonymous):

\[\int\limits_{400}^{\infty} kx^{-2}\] How do you do this to get an answer as 400?

OpenStudy (anonymous):

it's \[x \ge 400 \]otherwise..

razor99 (razor99):

srry i am trying to grt help.

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} kx ^{-2} =1\] . This does not seem to give the answer. Are you sure the question is typed correctly?

OpenStudy (anonymous):

It's 400 at the bottom... I changed it underneath the question

OpenStudy (anonymous):

ok then that integral is just k/400=1. k=400.

OpenStudy (anonymous):

But, why?

OpenStudy (anonymous):

\[\int\limits_{400}^{\infty}kx ^{-2}= -k/x\] put limits 1/infinity is 0 substitue 400 you'll get the answer.

OpenStudy (anonymous):

I didn't get that...

OpenStudy (anonymous):

integral of x^a is x^(a+1)/(a+1) do you know that?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

ok so now you know it :P :D

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

I did know it, sorry... So what then?

OpenStudy (anonymous):

Here a=-2 you get -1/x as the integral which on putting limits becomes 1/400.

OpenStudy (anonymous):

Ah Ok. I see now :) Thanks. And there's a B part... (b) find the probability that such a cartridge has a life of at least 500 hours

OpenStudy (anonymous):

so not put limits as 500 to infinity,

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

No problem....